OpenStudy (anonymous):

In a classroom, 2/3 of the students have brown hair. Furthermore, 1/2 have both brown hair and brown eyes. The probability that a brown-haired student also has brown eyes equals A. 1/3 B. 2/3 C. 3/4 D. 3/5

OpenStudy (anonymous):

1/2 * 2/3 =?

OpenStudy (anonymous):

0.24?

OpenStudy (anonymous):

no 1 2 -- * -- 2 3 multiply the numerators together, then the denominators to get a new fraction

OpenStudy (anonymous):

Ohhhh lol my slow self 2/6 with reduces to 1/3

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yup, 2/3 of the students have brown hair and 1/2 of them have brown eyes = 1/3